OpenStudy (anonymous):

The density of copper is 10.0 g/mL. What is the volume in milliliters of a 6284-g ingot of copper?

OpenStudy (rogue):

\[D = \frac {m}{V}\]\[10.0 g/mL = \frac {6284 g}{V}\]\[V = \frac {6284 g}{10.0 g/mL} = 628.4 mL\]